OpenStudy (kira_yamato):

A curve has the eqn y = 2sin^2x - 5cos x. Find in terms of p the new value of y as x changes from π/2 to (π/2 - p), where p is small.

OpenStudy (kira_yamato):

\[y=2\sin^2x-5 \cos x\]

OpenStudy (anonymous):

hi kira, are you in the rate of change in calculus?

OpenStudy (kira_yamato):

What about it?

OpenStudy (anonymous):

is the prob about the rate of change?

OpenStudy (kira_yamato):

I'm not sure... But to me it seems like it's related to limits...

OpenStudy (anonymous):

so your class is now in limits or rate of change

OpenStudy (kira_yamato):

This question is posed to me by my friend... I don't know his course though...

OpenStudy (anonymous):

ok this is just a derivative f(x) - f(c) f(pi/2) - f(pi/2 -p) f '(c)= lim ) -------- = lim ----------------- x->c x - c pi/2->(pi/2 -p) pi/2 - (pi/2 - p) =2 sin2(pi/2)-5cos(pi/2) - 2 sin2(pi/2 -p)-5cos(pi/2 -p) --------------------------------------------- pi/2 - (pi/2 - p) you need to use trig and some identities to arrive in the derivative of f(x) here or dy/dx = Dx [ 2(sin 2x)^2 ]- 5 cos x] = 2[4sin2x cos2x] + 5 sinx

OpenStudy (kira_yamato):

Thanks

OpenStudy (anonymous):

or maybe the prob may be looking only on increment or change only.... use the formula.... change in y = f(x) - f(c) or use change in y = f(x+h) - f(x) ,,if h=x-c and x=pi/2 , c=pi/2 -p h=pi/2 -(pi/2 -p)= p where h=p=small change in y = f(x+h) - f(x) change in y = f(x+p) - f(x) change in y = 2sin^2(x+p) - 5cos (x+p) - [ 2sin^2x - 5cos x]

OpenStudy (anonymous):

change in y = 2sin^2(pi/2+p) - 5cos (pi/2+p) - [ 2sin^2x - 5cos x] you need to use trig and some identities to arrive in the change in y or delta y